OpenStudy (anonymous):

4x+2y=12 what is the x intercpt

OpenStudy (anonymous):

what is the x intercpt

OpenStudy (anonymous):

do you know y?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok well x = 2 and y =-4 :)

OpenStudy (anonymous):

how did you get that

OpenStudy (anonymous):

wait i'm answering the wrong question oops here let me show you how you can do yours

OpenStudy (anonymous):

if you want to find the x intercept you have to make y = 0 but to find y you have to make the x = 0 so in thos case we put y = 0 so this will be our equation 4x+2(0)=12 4x + 0 = 12 - 0 - 0 4x = 12 /4 /4 x = 3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np can i get a medal and you become a fan so whenever i'm online you will know so i can help?